Colloquium

A Mathematics Inquiry Learning at Kutztown (MILK) Lecture

3:30 p.m. Thursday, September 30, 2010

Boehm Hall 260

The Moore Method: An Introduction

Dr. Padraig M^{c}Loughlin Associate Professor of Mathematics

Abstract

We begin by introducing the audience to the classic ’Moore Method’ and its originator, R. L. Moore (1882 - 1974). We present Moore in his own words (some snippets from the film, Challenge in the Classroom) and outline the main features of the method. We describe a ’Moore Method’ course (of which the speaker was a student more than a dozen times at several universities). A course taught in this manner oft has as a purpose encouraging the students to create, write, and present mathematically sound proofs of theorems.

We transition to a discussion of ’Modified Moore Methods’(M^{3}) which include the speaker’s own modified Moore method. The way a course is conducted varies from instructor to instructor, but the content of the course is usually presented in whole or in part by the students themselves. Instead of using a textbook, the students are given a list of definitions and theorems which they are to prove and present in class, leading them through the subject material. The Moore method or a modified Moore method typically limits the amount of material that a class is able to ’cover,’ but promotes a depth of understanding that, we opine, is missing when one listens to a lecture.

We discuss ’Inquiry-Based Learning’ (IBL) and how the Moore method and modified Moore methods are but types of inquiry-based learning techniques that can be used to successfully (hopefully) aid student learning. We close with a preview of the next speaker’s talk, Dr. Stan Yoshinobu (California Polytechnic University), who will speak on the 14^{th} of October in more depth and with greater clarity about inquiry-based learning.

3:00 p.m. refreshments served

3:30 p.m. talk begins

The Mathematics Inquiry Learning at Kutztown (MILK) Lecture is funded by the generous support of the Academy for Inquiry-Based Learning, The Educational Advancement Foundation, and Mr. Harry Lucas, Jr.

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul S. Ache III, Chair